what is the electric flux passing through a Gaussian surface that contains 2 protons and 1 electron?


1
Expert's answer
2022-03-24T11:00:33-0400

The Gauss's law says

"\phi=\oint EdA=\frac{q_{enc}}{\epsilon_0}"

"\phi=\frac{2e-e}{\epsilon_0}=\frac{e}{\epsilon_0}"

"\phi=\frac{1.6*10^{-19}}{8.85*10^{-12}}=1.8*10^{-8}\:\rm N\cdot m^2/C"